OpenStudy (debpriya):

the product of all the factors of a number N is equal to the tenth power of the number. which of the following cannot be the factor of N? (a. 90 b.200 c.210 d. 126)?

OpenStudy (anonymous):

Anything to the tenth power ends in zero. So cross D out.

OpenStudy (anonymous):

Oops. I read the question wrong. The answer is D. My bad.

OpenStudy (owen3):

5^10= 9765625 does not end in zero

OpenStudy (owen3):

3^10 does not end in zero either

OpenStudy (anonymous):

5 or 0

OpenStudy (owen3):

3^10= 59049 ends in 9

OpenStudy (debpriya):

the number of factors will be 20. that is the first step I know but I don't know how to do it further.

OpenStudy (anonymous):

I give up. You can take it from here.

OpenStudy (owen3):

The prime factorization of 90 is 2*3*3*5 So the factors of 90 are 2, 3,3,5 2*3 , 2*3, 2*5 , 3*3, 3* 5, 3* 5 2*3*3, 2*3*5, 3*3*5 2*3*3*5 delete repeats 2,3,5 2*3, 2*5 , 3*3,3* 5 2*3*3, 2*3*5, 3*3*5 2*3*3*5 The product of the factors of 90 is 2*3*5*(2*3)*(2*5)*(3*3)*(3* 5)*(2*3*3)*(2*3*5)*(3*3*5)*(2*3*3*5)= 531441000000 90^10= 34867844010000000000

OpenStudy (owen3):

we can rule out 90

OpenStudy (debpriya):

i saw one more solution where they found the number of factors of N. now there is a formula where the product of the factors is equal to N^m/2 where m is the number of factors .....now since in this question N^10 so the number of factors is 20. then they found the prime factorization of 20 which is 2*2*5. so they said that N can be written as a product of three prime factors. now since all the options except 210 have 3 prime factors we have to rule out 210 (it has 4 prime factors)...i did not get the logic in this ..but the answer is 210

OpenStudy (owen3):

I'm doing it out by brute force, and not getting that answer

OpenStudy (owen3):

factors of 210 2,3,5,7 (2*3), (2*5), (2*7),(3*5), (3*7),(5*7) (2*3*5), (2*3*7), (2*5*7), (3*5*7) (2*3*5*7) the product of all these is 3782285936100000000, which is not equal to 210^10 did i leave out a factor of 210 ?

ganeshie8 (ganeshie8):

Heyy

OpenStudy (debpriya):

nope you did not leave a factor of 210 answer is 210...i m just wondering if there is a shorter way to do this

OpenStudy (owen3):

I see, I misread the question. 210 is not N , it is a factor of N.

OpenStudy (owen3):

I am curious to find this number N .

OpenStudy (debpriya):

i think N can have a lot of values

OpenStudy (owen3):

the simplest case , a prime number, is impossible p = p^10 a number with two factors, a,b, a*b*(a*b) = (a*b)^10

ganeshie8 (ganeshie8):

Let \(d\) be a factor of \(N\), then we have \[n = d*\dfrac{n}{d}\tag{1}\] As \(d\) ranges over all the \(\tau(n)\) factors of \(N\), we get \(\tau(n)\) such equations. Multiplying them all together gives \[n^{\tau(n)} = \prod\limits_{d\mid n}d* \prod\limits_{\frac{n}{d}\mid n}\frac{n}{d}\tag{2}\]

ganeshie8 (ganeshie8):

But as \(d\) runs over the divisors of \(n\), so does \(\dfrac{n}{d}\); hence, \( \prod\limits_{d\mid n}d = \prod\limits_{\frac{n}{d}\mid n}\frac{n}{d}\)

ganeshie8 (ganeshie8):

so we get \[n^{\tau(n)} = \left(\prod\limits_{d\mid n}d\right)^2 \] which is same as \[n^{\tau(n)/2} = \prod\limits_{d\mid n}d \]

ganeshie8 (ganeshie8):

We are given that the product of divisors equals \(n^{10}\) : \[n^{\tau(n)/2} = \prod\limits_{d\mid n}d=n^{10}\]

ganeshie8 (ganeshie8):

so we endup solving \[\tau(n) = 2*10 = 20\]

ganeshie8 (ganeshie8):

Looks i have used \(n\) and \(N\) interchangebly...

ganeshie8 (ganeshie8):

In order to have \(20\) factors, \(n\) can only be of below forms : \(p^{1}q^1r^{4}\) \(p^3q^4\) \(p^1q^{11}\) \(p^{19}\)

ganeshie8 (ganeshie8):

Clearly \(n\) cannot have more than 3 different primes in its prime factorization

ganeshie8 (ganeshie8):

so none of the factors of \(n\) are allowed to have more than \(3\) different prime factors

ganeshie8 (ganeshie8):

In order to have \(20\) factors, \(n\) can only be of below forms : \(p^{1}q^1r^{4}\) \(p^3q^4\) \(p^1q^{\color{red}{9}}\) \(p^{19}\)

OpenStudy (owen3):

are you including 1 as a factor

ganeshie8 (ganeshie8):

yes all positive factors

ganeshie8 (ganeshie8):

Btw, above work makes sense only if you know how to find the number of divisors of a given integer

ganeshie8 (ganeshie8):

Let \[n = {p_1}^{e_1}{p_2}^{e_2}{p_3}^{e_3}\cdots {p_r}^{e_r}\] Do you know how to find the value of number of positive divisors \(\tau\) of \(n\) ?

ganeshie8 (ganeshie8):

@debpriya

OpenStudy (owen3):

Supplemental question. Find the smallest number N such that the product of all the factors of N is equal to the tenth power of the number.

OpenStudy (owen3):

A couple of comments. \( \large \prod\limits_{\frac{n}{d}\mid n}\frac{n}{d}\) is the same as \(\prod\limits_{d\mid n}d \), by listing the divisors backwards. And you proved an interesting result. $$n^{\tau(n)/2} = \prod\limits_{d\mid n}d$$ http://primes.utm.edu/glossary/xpage/tau.html

OpenStudy (debpriya):

@ganeshie8 i understood it. Thank you so much :)

ganeshie8 (ganeshie8):

Np :) see if you can find the smallest N whose product of divisors equals the 10th power of N